OpenStudy (anonymous):

find the equation of the tangent line to the graph of the function at the given point y=arctan 2/x, (2, pi/4)

OpenStudy (anonymous):

well just differentiate it then find the gradient m since the x value is given then y -y_1 = m(x-x_1) i.e eqn. of tangent line

OpenStudy (anonymous):

\[f(x)=\arctan(\frac{2}{x})\] \[f'(x)=\frac{-2}{x^2+4}\] \[f'(2)=-\frac{1}{4}\] then point - slope

OpenStudy (amistre64):

or... :)\[tan(y)=\frac{2}{x}\] \[sec^2(y)y'=-\frac{2}{x^2}\] \[y'=-\frac{2cos^2(y)}{x^2}\]

OpenStudy (amistre64):

\[y'=-\frac{2~cos^2(45^o)}{4}\] \[y'=-\frac{cos^2(45^o)}{2}\] \[y'=-\frac{1}{(\sqrt{2})^22}\] \[y'=-\frac{1}{4}\] lol